can someonnnneeeeeee help plsssss ;(?
solve this equation. check your answer.
3x-1/x-2 = 5/x-2 +1

Answer by Alan3354(69443)

3x-1/x-2 = 5/x-2 +1
---------
I think you mean
(3x-1)/(x-2) = 5/(x-2) +1
Multiply by x-2
3x-1 = 5 + x-2 = 3+x
2x = 4
x = 2
----------
x = 2 makes the denomination zero, so no solution